# Canonical Representation and Force-Based Pretraining of 3D Tactile for Dexterous Visuo-Tactile Policy Learning

### v1

@@ I INTRODUCTION
Human hands play a crucial role in daily life [1], enabling a wide range of tasks such as opening boxes and flipping objects. This level of dexterity is essential for integrating robots into everyday human activities. Visual base imitation learning has shown great potential in teaching dexterous hands to perform various tasks [2, 3, 4]. While simpler tasks like pick-and-place operations can achieve high success rates, more fine-grained and contact-rich tasks—such as in-hand manipulation, remain significantly challenging. These tasks involve precise control of force, nuanced coordination of different fingers, and continuous feedback during manipulation. A key factor in successfully executing such tasks is tactile sensing [5].
To enable dexterous hands to perceive contact, current approaches typically equip them with tactile sensors. These sensors can be mainly categorized into vision-based tactile sensors, such as GelSight [6], DIGIT [7], and distributed tactile sensors, like uSkin [8]. Distributed tactile sensors are particularly well-suited for various robotic structures due to their small size, which allows for easy integration. Their robustness also makes them reliable in diverse environments, leading to widespread use in many systems [9, 10, 11].
However, distributed tactile sensors typically cover a large area of the dexterous hand [12], resulting in high-dimensional input. Additionally, different dexterous hands often use different types of distributed tactile sensors with varying sensor distributions, leading to a lack of large-scale standardized datasets. This poses unique challenges in learning tactile features effectively to improve dexterous manipulation.
Considering the power of visual backbones, many works [13, 4] convert tactile data into 2D images to reduce the complexity of learning useful tactile features. While this transformation aids feature learning, it leads to the change and loss of part spatial information between different taxels. To preserve these spatial relationships, most approaches represent 3D tactile data as a graph and use graph neural networks (GNNs)[14] to encode tactile signals[12, 15]. However, these methods primarily focus on specific tasks and require a large amount of data to learn effective features. Inspired by the success of pretraining strategy in vision-based learning, T-DEX [4] collects tactile play data through self-supervised interaction with various objects. This pretraining improves the efficiency of feature learning and enhances diverse downstream robotic manipulation tasks. However, it still relies on 2D images as the tactile representation. As a result, efficiently learning 3D tactile data features for dexterous manipulation remains a challenge.
@@
To address the difficulty of 3D tactile feature learning, we first propose a novel canonical representation of 3D tactile data, which canonicalizes the coordinates of taxels in each sensor into a unified coordinate system. This canonicalization aligns the features of differently distributed sensors and reduces the feature space. Additionally, it amplifies the distances between taxels within the same sensor, facilitating the capture of more localized features. We further propose a force-based, self-supervised prediction task for pretraining 3D tactile data, given the importance of force usage in object manipulation. The pretraining tasks include both masked local force prediction and net force prediction, encouraging the encoder to learn features related to both local and net force relationships.
To demonstrate the effectiveness of our method, we incorporate the pretrained tactile encoder into an imitation learning framework and evaluate its manipulation performance in the real world on four fine-grained, contact-rich tasks: open box, reorientation, flip, and assembly. Comparative results demonstrate the effectiveness of our method compared to other baselines. Ablation studies confirm the importance of our proposed canonical representation and force-based pretraining. Additionally, our analysis shows that the policy effectively utilizes both the spatial and force information from the 3D tactile data.
@@ II-A Tactile for Dexterous Manipulation
Tactile sensing has been widely used to enhance dexterous manipulation. Current approaches primarily utilize either vision-based tactile sensors [6] or distributed tactile sensors [8]. Due to their larger size, vision-based tactile sensors are typically only mounted on the fingertips of dexterous hands [16]. In contrast, distributed tactile sensors can cover a larger area [10]. Thus, we choose to use distributed tactile sensors in this work.
@@ II-B Tactile Representaiton
Different tactile representations convey various types of information and can be encoded using different methods. For low-dimensional tactile data, directly applying MLP to flattened tactile readings [20] or converted 3D vectors [21] can still capture useful tactile features. However, as the tactile sensor coverage increases, the dimensionality of the data grows significantly, making direct encoding of tactile readings inefficient. To leverage powerful visual backbones, some methods convert raw tactile readings into RGB images by mapping the tri-axis forces to three channels [22, 13, 23], using visual backbones for encoding. However, such 2D information changes the inherent spatial relationship of taxels in the same sensor and does not contain the spatial relationship of taxels in different sensors that are distributed on the different parts of the robot. To preserve spatial relationships, graph-based methods have been applied to tactile data by treating each taxel as a node, connecting them with either predefined [15] or dynamically changing graph [12]. Nevertheless, the representation in these works only use a subset of the available tactile information in 3D space, such as the 3D position [12] or 3D force of the taxels [15]. Our work fully leverages both the 6D pose and 3D force of each taxel, and we propose a novel canonical representation to more effectively learn features from such complex tactile data.
@@ III ROBOT SYSTEM SETUP
As shown in Fig. 1, our system consists of a 6-Dof JAKA MiniCobo robot arm and a 16-Dof Leap Hand [29] dexterous hand with four fingers. The Leap Hand is equipped with PaXini tactile sensors, Each finger has two type of sensors: one for the fingertip and another for the fingerpad. Both types of sensors have a 3x5 array of taxels, but taxel distribution is slightly different. Each taxel measuring tri-axial forces $\mathbf{F}\in\mathbb{R}^{3}$. A single Intel RealSense D415 camera is mounted diagonally on the robot to capture visual information.
@@ IV METHOD
We focus on the problem of leveraging 3D tactile data from distributed tactile sensors for learning visuo-tactile dexterous manipulation policies. To reduce the difficulty of learning features from complex 3D tactile data, we canonicalize the data into a unit coordinate system, as described in IV-A. We then pretrain the tactile encoder using self-supervised force-based prediction tasks to enhance local and net force feature learning, detailed in IV-B. This pretrained tactile encoder is subsequently used for visuo-tactile policy learning, as outlined in IV-C.
@@ IV-A Canonical Tactile Representation
To preserve the spatial relationships of each taxel, we aim to use 3D tactile data instead of converting it into a 2D image. For each taxel of the distributed tactile sensor, in addition to the 3D force $\mathbf{F}$, we can also obtain the 6D pose $\mathbf{P}\in\mathbb{R}^{6}$ by computing forward kinematics. This information shows how the force is applied at every frame. However, since a large number of taxels are distributed across different parts of the fingers, using this 9D tactile representation results in a vast feature space, making it difficult to learn meaningful tactile features. Additionally, though the taxels within a sensor are distributed sparsely, the distances between them are very small (e.g., less than 2 millimeters), making it challenging to capture local features within the same sensor.
To address the challenges, we propose to canonicalize the 9D tactile representation. Specifically, we normalize each taxel’s coordinate within the same sensor into a unit coordinate system (ranging from -1 to 1 for each axis) by computing the diagonal length of the original coordinates within the sensor’s coordinate system. As shown in Fig. 3, the 3D position of each taxel in the unit coordinate system is denoted as $\mathbf{T}\in\mathbb{R}^{3}$. However, this representation only captures the spatial relationships between taxels within the same sensor, without accounting for the spatial relationships between different sensors. Therefore, we also include the 6D pose of each sensor’s origin with respect to the hand’s base, denoted as $\mathbf{P}^{s}$, into the representation. As a result, the representation for each taxel is represented as $\mathbf{R}=[\mathbf{P}^{s},\mathbf{T},\mathbf{F}]$. Although this representation has a higher dimension, it effectively reduces the feature space because the features of different sensors become more aligned due to the canonicalized coordinates. Additionally, this canonicalization amplifies the relative distance between taxels within the same sensor, making their features more distinguishable for the neural network. This facilitates the capture of more localized features for each taxel.
@@ IV-C Visuo-Tactile Policy Learning
After obtaining the pretrained tactile encoder, we utilize an imitation learning framework to learn visuo-tactile policy for dexterous manipulation tasks. Specifically, we adopt the diffusion policy [33] as our backbone. However, we replace the vision backbone from ResNet [34] to DinoV2 [35] for improved visual feature extraction. Unlike the original diffusion policy, we incorporate tactile data as an additional input, using the pretrained tactile encoder to encode this data. The output features are then concatenated with visual features as a condition for policy learning. We fine-tune the encoder during downstream tasks training, following diffusion policy [33]. In addition to visual and tactile inputs, we include the robot’s proprioceptive state, including 3D position, 4D quaternion of the robot arm, and 16D joint position of the dexterous hand. Notably, our action space consists of target actions that humans intended to achieve during teleoperation, rather than states, as these actions implicitly capture the use of force, which is crucial for accomplishing a variety of tasks [36, 37].
@@ V EXPERIMENTS
Figure 4: Visualization of Our Policy’s Rollout on Four Fine-Grained, Contact-Rich Tasks. Note this view is the view of the robot’s observation.
@@ V-A Dexteous Manipulation Tasks
We conduct experiments on four dexterous fine-grained, contact-rich manipulation tasks, as shown in Fig. 4. Each experiment run will be limited to a maximum of 600 steps. Each method will be evaluated on each task of 10 experiment runs.
1) Open Box
This task requires the robot to open a box using the thumb and index finger. The robot needs to first reach the box, grasp the upper part, and then carefully adjust its finger to open the box without pushing it. The challenge is maintaining a firm hold on the upper part during opening to prevent it from loosening and falling. The box is placed in a random pose within an 18x12 cm area for each run. Success is achieved if the upper part of the box stays in place after opening.
2) Reorientation
This task requires the robot to continuously reorient a bottle until it points in a specific direction. The robot needs to reach the bottle and coordinate its four fingers to reorient it without pushing it down. The challenge is the precise coordination of the fingers, and the task is long-horizon. The bottle is placed in a random pose within an 18x12 cm area for each run. Success is achieved if the bottle is within 10 degrees of the target direction.
3) Flip
This task requires the robot to flip a bottle cap using the thumb, middle, and index finger. The robot needs to reach the cap, grasp it, lift one side of the cap, and use the index finger to flip the cap. The challenge involves precise finger coordination and force application, with severe occlusion and ambiguity during the process. The cap is placed in a fixed position with random orientations, and success is achieved if the cap is flipped by 180 degrees.
4) Assembly
This task requires the robot to grasp one part of a box and assemble it with another. The robot needs to reach, grasp, move, and gradually insert one part into the other. The challenge is making fine adjustments based on feedback while handling high occlusion and ambiguity. The box parts are in fixed positions, and success is achieved when one part is successfully inserted into the other.
@@ V-B Baselines
We compare our method with the following baselines, which all use the same visual backbone, diffusion policy backbone, visual observation, robot proprioceptive state, and action space as ours, but with different types of tactile representation and pertaining. 1) DP: We implement the diffusion policy without using tactile data for this baseline. 2) HATO: HATO [20] uses MLP to directly encode the tactile readings. We flatten 360-dimensional force values and use MLP to encode tactile data for this baseline. 3) T-DEX: T-DEX [4] convert the raw tactile readings into 2D image, and pretrain with BYOL [27]. For this baseline, we follow their procedure, first converting raw, tactile readings into 2D images, then pretrain on our own collected dataset and we use the pretrained encoder for diffusion policy learning. 4) GNN: We use the 9D tactile representation (e.g., the 6D pose of each taxel with respect to the hand’s base with 3D tactile force values) as input for this baseline, use graph attention networks [38] to encode tactile, which is the same GNN backbone as ours. We do not pretrain for this baseline.
@@ V-C Manipulation Policy Comparsion
As shown in Tab. I, our approach achieves the highest success rate across all tasks. While most of the baselines perform well on the open box and reorientation tasks, they struggle with the assembly and flip tasks. Interestingly, we found that even without tactile feedback, DP still achieves a high success rate on open box and reorientation tasks. This is mainly because these tasks do not involve significant occlusion or ambiguity during manipulation, allowing DP to successfully find and manipulate objects using visual input and robot state alone. In contrast, even with rich tactile information, GNN consistently fails across all tasks, we observe that the finger or the hand usually shakes during the manipulation, preventing it from finishing the task. Compared to GNN, HATO, which only uses tactile force values, is able to accomplish some tasks, demonstrating the difficulty of learning spatial and force information from 3D tactile data simultaneously. T-DEX performs better than the other baselines, showing that even 2D tactile data with pretraining can lead to high success rates, though it struggles with the flip task.
@@ V-D Importance of Representation and Pretraining
To validate the effectiveness of our proposed canonical representation and force-based pretraining, we conduct ablation studies across all tasks. As shown in Tab. II, using the canonical representation of tactile data, even without pretraining, the trained policy achieves a 48% success rate. This indicates that the representation significantly reduces the complexity of learning from 3D tactile data. With force-based pretraining, the success rate increases to 78%, demonstrating that pretraining indeed enhances the performance of the visuo-tactile policy. By combining both canonical representation and pretraining, Ours can achieve robust success rate across both easy and difficult tasks.
@@ V-E Effect of Force-based Pretraining Tasks
To validate the effectiveness of our pretraining tasks, we conducted experiments using only one pretraining task at a time. As shown in Tab. III, omitting either pertaining task leads the encoder to focus solely on either local or net force features, failing to perform the manipulation effectively. The results also show that net force prediction is more critical for achieving the manipulation tasks.

### v2

@@ I INTRODUCTION
Human hands are vital in daily life [1], enabling a wide range of tasks such as opening boxes and flipping objects. This level of dexterity is essential for integrating robots into everyday human activities. Vision-based imitation learning has shown great potential in teaching dexterous hands to perform various tasks [2, 3, 4]. While simpler tasks like pick-and-place operations can achieve high success rates, more fine-grained and contact-rich tasks—such as flip object, remain significantly challenging. These tasks involve precise control of force, nuanced coordination of different fingers, and continuous feedback during manipulation. A key factor in successfully executing such tasks is tactile sensing [5].
To enable dexterous hands to perceive contact, current approaches typically equip them with tactile sensors. These sensors can be mainly categorized into vision-based tactile sensors, such as GelSight [6], DIGIT [7], and distributed tactile sensors, like uSkin [8]. Distributed tactile sensors are particularly well-suited for various robotic structures due to their small size, which allows for easy integration. Their robustness also makes them reliable in diverse environments, leading to widespread use in many systems [9, 10, 11]. However, distributed tactile sensors typically have lots of taxels and cover large areas on dexterous hand [12], leading to high-dimensional input. Moreover, different dexterous hands often use different types of distributed tactile sensors with varying sensor distributions, resulting in a lack of large-scale standardized datasets. This poses challenges in effectively learning tactile features for dexterous manipulation.
Considering the power of visual backbones, many works [13, 4] convert tactile data into 2D images to reduce the complexity of learning useful tactile features. However, this transformation leads to the change and loss of part spatial information between different taxels. To preserve these spatial relationships, most approaches represent 3D tactile data as a graph and use graph neural networks (GNNs)[14] to encode tactile signals[12, 15]. However, these methods primarily focus on specific tasks and require a large amount of data to learn effective features. Inspired by the success of the pretraining strategy in vision-based learning, T-DEX [4] collects tactile play data through interaction with various objects and pretrain tactile encoder with self-supervised learning. This pretraining improves the efficiency of feature learning and enhances diverse downstream robotic manipulation tasks. However, it still relies on 2D images as the tactile representation. As a result, efficiently learning 3D tactile data features for dexterous manipulation remains a challenge.
@@
To address the difficulty of 3D tactile feature learning, we first propose a novel canonical representation of 3D tactile data, which canonicalizes the coordinates of taxels in each sensor into a unified frame. This canonicalization aligns the features of differently distributed sensors and reduces the feature space. Additionally, it amplifies the distances between taxels within the same sensor, facilitating the capture of more localized features. We further propose a force-based, self-supervised prediction task for pretraining 3D tactile data, given the importance of force usage in object manipulation. The pretraining tasks include both masked local force prediction and net force prediction, encouraging the encoder to learn features related to both local and net force relationships.
To demonstrate the effectiveness of our method, we integrate the pretrained tactile encoder into an imitation learning framework and evaluate it in the real world on four fine-grained, contact-rich tasks: open box, reorientation, flip, and assembly. Comparative results demonstrate the effectiveness of our method compared to other baselines. Ablation studies confirm the importance of our proposed canonical representation and force-based pretraining. Additionally, our analysis shows that the policy effectively utilizes both the spatial and force information from the 3D tactile data.
@@ II-A Tactile for Dexterous Manipulation
Tactile sensing has been widely used to enhance dexterous manipulation. Current approaches primarily utilize either vision-based tactile sensors [6] or distributed tactile sensors [8]. Vision-based tactile sensors are typically only mounted on the fingertips of dexterous hands [16], due to their large size. In contrast, distributed tactile sensors can cover a larger area [10]. Thus, we choose to use distributed tactile sensors in this work.
@@ II-B Tactile Representaiton
Different tactile representations convey various types of information and can be encoded using different methods. For low-dimensional tactile data, directly applying MLP to flattened tactile readings [20] or converted 3D vectors [21] can capture useful tactile features. However, as sensor coverage increases, the dimensionality of the data grows significantly, making direct encoding of tactile readings inefficient. To leverage powerful visual backbones, some methods convert raw tactile readings into RGB images by mapping the tri-axis forces to three channels [22, 13, 23], using visual backbones for encoding. However, such 2D information changes the inherent spatial relationship of taxels in the same sensor and does not contain the spatial relationship of taxels in different sensors that are distributed on the different parts of the robot. To preserve spatial relationships, graph-based methods have been applied to tactile data by treating each taxel as a node, connecting them with either predefined [15] or dynamically changing graph [12]. Nevertheless, the representation in these works only use a subset of the available tactile information in 3D space, such as the 3D position [12] or 3D force of the taxels [15]. Our work fully leverages both the 6D pose and 3D force of each taxel, and we propose a novel canonical representation to more effectively learn features from such complex tactile data.
@@ III ROBOT SYSTEM SETUP
As shown in Fig. 1, our system consists of a 6-Dof JAKA MiniCobo robot arm and a 16-Dof Leap Hand [29] dexterous hand with four fingers. The Leap Hand is equipped with PaXini tactile sensors, Each finger has two types of sensors: one for the fingertip and another for the fingerpad. Both types of sensors have a 3x5 array of taxels, but taxel distribution is slightly different. Each taxel measuring tri-axial forces $\mathbf{F}\in\mathbb{R}^{3}$. A single Intel RealSense D415 camera is mounted diagonally of the robot to capture visual information.
@@ IV METHOD
We focus on the problem of leveraging 3D tactile data from distributed tactile sensors for learning visuo-tactile dexterous manipulation policies. To reduce the difficulty of learning features from complex 3D tactile data, we canonicalize the data into a unit frame, as described in IV-A. We then pretrain the tactile encoder using self-supervised force-based prediction tasks to enhance local and net force feature learning, detailed in IV-B. This pretrained tactile encoder is subsequently used for visuo-tactile policy learning, as outlined in IV-C.
@@ IV-A Canonical Tactile Representation
To preserve the spatial relationships of each taxel, we aim to use 3D tactile data instead of converting it into a 2D image. For each taxel of the distributed tactile sensor, in addition to the 3D force $\mathbf{F}$, we can also obtain the 6D pose $\mathbf{P}\in\mathbb{R}^{6}$ by computing forward kinematics. This information shows how the force is applied at every step. However, since a large number of taxels are distributed across different parts of the fingers, using this 9D tactile representation results in a vast feature space, making it difficult to learn meaningful tactile features. Additionally, though the taxels within a sensor are distributed sparsely, the distances between them are very small (e.g., less than 4 millimeters), making it challenging to capture local features within the same sensor.
To address the challenges, we propose to canonicalize the 9D tactile representation. Specifically, we normalize each taxel’s coordinate within the same sensor into a unit frame (ranging from -1 to 1 for each axis) by computing the diagonal length of the original coordinates within the sensor frame. As shown in Fig. 3, the 3D position of each taxel in the unit frame is denoted as $\mathbf{T}\in\mathbb{R}^{3}$. However, this representation only captures the spatial relationships between taxels within the same sensor, without accounting for the spatial relationships between different sensors. Therefore, we also include the 6D pose of each sensor’s origin with respect to the hand’s base, denoted as $\mathbf{P}^{s}$, into the representation. As a result, the representation for each taxel is represented as $\mathbf{R}=[\mathbf{P}^{s},\mathbf{T},\mathbf{F}]$. Although this representation has a higher dimension, it effectively reduces the feature space because the features of different sensors become more aligned due to the canonicalized coordinates. Additionally, this canonicalization amplifies the relative distance between taxels within the same sensor, making their features more distinguishable for the neural network. This facilitates the capture of more localized features for each taxel.
@@ IV-C Visuo-Tactile Policy Learning
After pretraining the tactile encoder, we use imitation learning to learn visuo-tactile policy for dexterous manipulation. Given the diffusion model’s ability to model complex action distributions [33], we adopt the diffusion policy [34] as our backbone. We replace the vision backbone with DinoV2 [35] for better visual feature extraction and integrate tactile data as an additional input, encoded by the pretrained tactile encoder. The tactile features are concatenated with visual features for policy learning. We fine-tune the encoder during downstream tasks training, following diffusion policy [34]. We also include the robot’s proprioceptive state, including 3D position, 4D quaternion of the arm, and 16D joint position of the hand. Our action space is target actions rather than states, as these actions implicitly capture force usage, crucial for accomplishing diverse tasks [36, 37].
@@ V EXPERIMENTS
Figure 4: Visualization of Our Policy’s Rollout on Four Fine-Grained, Contact-Rich Tasks. Note this is the view of the robot’s observation.
@@ V-A Dexteous Manipulation Tasks
We conduct experiments on four dexterous fine-grained, contact-rich manipulation tasks, as shown in Fig. 4. Each experiment run will be limited to a maximum of 600 steps. Each method will be evaluated on each task of 10 experiment runs. 1) Open Box: This task requires the robot to open a box using the thumb and index finger. The robot needs to first reach the box, grasp the upper part, and then carefully adjust its finger to open the box without pushing it. The challenge is maintaining a firm hold on the upper part during opening to prevent it from loosening and falling. The box is placed randomly within an 18x12 cm area for each run. Success is achieved if the upper part of the box stays in place after opening. 2) Reorientation: This task requires the robot to continuously reorient a bottle until it points in a specific direction. The robot needs to reach the bottle and coordinate its four fingers to reorient it without pushing it down. The challenge is the precise coordination of the fingers, and the task is long-horizon. The bottle is placed in a random pose within an 18x12 cm area for each run. Success is achieved if the bottle is within 10 degrees of the target direction. 3) Flip: This task requires the robot to flip a bottle cap using the thumb, middle, and index finger. The robot needs to reach the cap, grasp it, lift one side of the cap, and use the index finger to flip the cap. The challenge involves precise finger coordination and force application, with severe occlusion and ambiguity during the process. The cap is placed in a fixed position with random orientations, and success is achieved if the cap is flipped by 180 degrees. 4) Assembly: This task requires the robot to grasp one part of a box and assemble it with another. The robot needs to reach, grasp, move, and gradually insert one part into the other. The challenge is making fine adjustments based on feedback while handling high occlusion and ambiguity. The box parts are in fixed positions, and success is achieved when one part is successfully inserted into the other.
@@ V-B Baselines
We compare our method with the following baselines, which all use the same visual backbone, diffusion policy backbone, visual observation, robot proprioceptive state, and action space as ours, but with different types of tactile representation and pertaining. 1) DP: We implement the diffusion policy without using tactile data or pertaining for this baseline. 2) HATO: HATO [20] uses MLP to directly encode the tactile readings. We flatten 360-dimensional force values and use MLP to encode tactile data for this baseline. 3) T-DEX: T-DEX [4] convert the raw tactile readings into 2D image, and pretrain with BYOL [27]. For this baseline, we follow their procedure, first converting raw, tactile readings into 2D images, then pretrain on our own collected dataset and we use the pretrained encoder for diffusion policy learning. 4) GNN: We use the 9D tactile representation (e.g., the 6D pose of each taxel with respect to the hand’s base with 3D tactile force values) as input for this baseline, use graph attention networks [38] to encode tactile, which is the same GNN backbone as ours. Since there are no pertaining strategy designed especially for 9D tactile representation, We do not pretrain for this baseline.
@@ V-C Manipulation Policy Comparsion
As shown in Tab. I, our approach achieves the highest success rate across all tasks. Most baselines perform well on the open box and reorientation tasks but struggle with the assembly and flip tasks. Interestingly, we found that even without tactile feedback, DP still achieves a high success rate on open box and reorientation tasks. This is mainly because these tasks do not involve significant occlusion or ambiguity during manipulation, allowing DP to successfully find and manipulate objects using visual input and robot state alone. In contrast, even with rich tactile information, GNN consistently fails across all tasks, we observe that the finger or the hand usually shakes during the manipulation, preventing it from finishing the task. Compared to GNN, HATO, which only uses tactile force values, is able to accomplish some tasks, demonstrating the difficulty of learning spatial and force information from 3D tactile data simultaneously. T-DEX performs better than the other baselines, showing that even 2D tactile data with pretraining can achieve high success rates, though it struggles with the flip task.
@@ V-D Importance of Representation and Pretraining
To validate the effectiveness of our canonical representation and force-based pretraining, we conduct ablation studies across all tasks. As shown in Table II, using the canonical representation achieves a 48% success rate, even without pretraining. However, without the canonical representation, even with pretraining, the policy fails to complete any task. We observed that the policy directly moves to a specific hand joint position upon starting inference and then repeats similar actions. Analyzing the output of tactile encoder, we found that without the canonical representation, the encoder outputs similar features for taxels within the same sensor pad, failing to perceive fine-grained differences. This validates the necessity of the canonical representation. Based on such representation, pretraining further enables the encoder to learn more useful features, increasing the success rate to 78%.
@@ V-E Effect of Force-based Pretraining Tasks
To validate the effectiveness of our pretraining tasks, we conducted experiments using only one pretraining task at a time. As shown in Tab. III, omitting either pertaining task leads the encoder to focus solely on either local or net force features, resulting in a significant performance drop in all tasks. The results also show that net force prediction is more critical for achieving the tasks.